OpenStudy (anonymous):

Please help me? Been asking this question for days and no one has helped at all. :\

OpenStudy (anonymous):

So i have to fill out this table and im not sure exactly how to do it. I think i have some sort of understanding, but im still confused. Heres the first part, what i need to do. Adjustable Rate Mortgage: 3% with terms 5/1 with a 2/6 cap for 30 years (Assume the interest rate increases by 1.25% after the initial period and every 10 years thereafter.) I have to find the Monthly Payment, the # of Payments, and the Total Cost for Each Period for these years. 1-5 6-15 16-25 26-30 And then the Total.

OpenStudy (anonymous):

I think i have the percentages down. So i figured out the 3% with terms 5/1 stuff and figured out the percentages for each year. Year 1 3% Year 2 3% Year 3 3% Year 4 3% Year 5 3% Year 6 5% Year 7 7% Year 8 9% Year 9 11% Year 10 13% Year 11 15% Year 12 17% Year 13 19% Year 14 21% Year 15 23% Year 16 25% Year 17 27% Year 18 29% Year 19 31% Year 20 33% Year 21 35% Year 22 37% Year 23 39% Year 24 41% Year 25 43% Year 26 45% Year 27 47% Year 28 49% Year 29 51% Year 30 53%

jimthompson5910 (jim_thompson5910):

If the interest rate reaches 53%, then that's very extreme and very illegal. You'd probably never pay off the mortgage at that point since much of the payment (maybe all of it) would be interest alone.

OpenStudy (anonymous):

Oh ok cause i just took the 5% and just kept adding 2%. So maybe that was where i went wrong?

jimthompson5910 (jim_thompson5910):

that's where the " 2/6 cap" comes in. I'm currently looking up to see whether the "6" means "6% is the highest the interest rate can go" OR if it means "6% is the highest INCREASE allowed". If that is the case, then the highest the interest rate is 3+6 = 9%

jimthompson5910 (jim_thompson5910):

ok this page says it's the second interpretation http://qna.mortgagenewsdaily.com/questions/arm-caps-initial-periodic-and-lifetime

jimthompson5910 (jim_thompson5910):

so the lowest interest rate is 3% the highest interest rate is 9% the interest rate cannot go over 9% due to the cap

OpenStudy (anonymous):

Ok so i have a formula here for how to find the monthly payment. Its \[M = B \frac{ i(1+i)^nt}{ (1+i)^nt - 1 }\] B = the balance which is $1,149,000 i = Periodic Interest rate which is 3%? n = Number of Payments per Year which is ? t = Number of years remaining on the loan which is 30 or after the 5 years, 25?

OpenStudy (anonymous):

The t is supposed to be part of the exponent. Supposed to be ^nt. @jim_thompson5910

OpenStudy (anonymous):

Oh! Would the number of monthly payments per year be 12? Since its paid monthly? Lol Ok i got that. What about t? Am i right on that?

jimthompson5910 (jim_thompson5910):

it's a 30 year mortgage, so t = 30 the monthly payment stays fixed throughout the entire life of the mortgage. This is assuming you do NOT refinance

jimthompson5910 (jim_thompson5910):

n = 12 because you pay monthly (12 months = 1 yr) so n*t = 12*30 = 360 months ie you make 360 payments of whatever the monthly payment is to pay back the principal and pay interest as well

OpenStudy (anonymous):

Ok so heres my work... (Sorry, might be a little hard to read) (I got 0.0025 by taking the initial interest rate, 3%, and dividing 0.03 by 12, monthly payments, and got that.) = 1,149,000[0.0025(1+0.0025)^360/(1+0.0025)^360 - 1] = 1,149,000[0.0025(1.0025)^360/(1.0025)^360 - 1] = 1,149,000[0.0025(2.4568)/(2.4568) - 1] = 1,149,000(0.0042) = $4,825.80

OpenStudy (anonymous):

Oops. I think i did that wrong. Wolframalpha got $4,844.23.

OpenStudy (anonymous):

Im still confused. Are you still going to help? Its getting late and i need to go soon. :\ @jim_thompson5910

OpenStudy (anonymous):

@tkhunny Are you able to help?

OpenStudy (tkhunny):

One piece at a time. 1) Ignore EVERYTHING except the 3%. Calculate the remaining balance at the end of the 5 years. 2) Calculate a new payment with 25 years left, given the next year's interest rate. Pay that for one year and recalculate the remaining balance after that year. 3) Calculate a new payment with 24 years left, given the next year's interest rate. Pay that for one year and recalculate the remaining balance after that year. ... 4) Thoroughly agreeing with Mr. Thompson, do this one last time when you get to 9%. You should be done after that.

OpenStudy (anonymous):

So would i do this to find out what the monthly payment would be? (Sorry for all the questions, i suck at math and im still confused.) M = 1,149,000[0.03(1+0.03)^60/(1+0.03)^60 – 1]

OpenStudy (anonymous):

Look at this website. Does this look right to you? http://www.zillow.com/mortgage/calculator/payment/ModernPaymentCalculatorAdvancedReportPage.htm? {"homePrice":1149000,"downPayment":0,"term":360,"arm":false,"rate":3,"includeTaxesInsurance":true,"propertyTaxRate":1.2,"annualHomeownersInsurance":800,"includePMI":true,"monthlyHOA":0} @jim_thompson5910

OpenStudy (anonymous):

@ganeshie8 @mathstudent55 @Directrix Anyone?!?! ^^^^^

jimthompson5910 (jim_thompson5910):

I'm trying out the zillow calculator and it's stuck on "Loading..."

jimthompson5910 (jim_thompson5910):

can you take a screenshot of what you see?

OpenStudy (anonymous):

Its too long. :\ Copy and paste the entire thing into your search bar. Worked for me.

jimthompson5910 (jim_thompson5910):

oh the link cut off, i see

OpenStudy (anonymous):

Yeah, sorry. :\

jimthompson5910 (jim_thompson5910):

ok that works and looks good. BUT it only applies for years 1-5 since the interest rate is 3% on year 6, you'll have to bump the interest up to 5% (until you reach year 7) and so on

OpenStudy (anonymous):

Oh! Ok so when i hit 9%, stop?

jimthompson5910 (jim_thompson5910):

yeah once you hit 9%, you don't increase the interest rate any higher of course, that doesn't mean you stop making payments once you reach 9% (you still have 30 years)

OpenStudy (anonymous):

Oh ok. So it just stays at 9% then?

jimthompson5910 (jim_thompson5910):

yeah the remaining bunch of years will have the interest rate fixed at 9%

OpenStudy (anonymous):

Omw. Thank you!!! Sorry, one more quick question, for the column that asks me to give the total cost for each period, do i take the remaining balance from the first month and subtract the 60th month from that?

jimthompson5910 (jim_thompson5910):

Simple example: say someone loaned you $100 and asked for $110 back the total cost of that loan (to you) is $110. You pay back the initial principal of $100 and interest of $10

jimthompson5910 (jim_thompson5910):

if you split up the payments into pieces, say 11 periods, then 110/11 = 10 which means the monthly payment is $10 a month for 11 months each month, the total cost will go up by $10 because the total cost is cumulative

OpenStudy (anonymous):

Oh ok. Alright thank you soo much! I was so stuck on this. :)

jimthompson5910 (jim_thompson5910):

you're welcome